Exercise 11.2

Ajax Cleaning Products is a medium-sized firm operating in an industry dominated by one large firm—Tile King. Ajax produces a multiheaded tunnel wall scrubber that is similar to a model produced by Tile King. Ajax decides to charge the same price as Tile King to avoid the possibility of a price war. The price charged by Tile King is $20,000. Ajax has the following short-run cost curve:

TC=800,000−5,000Q+25Q2TC=800,000−5,000Q+25Q2

What is the marginal cost (MC) curve for Ajax?

−5,000+25Q−5,000+25Q

800,000Q−5,000+25Q800,000Q−5,000+25Q

−5,000Q+50Q−5,000Q+50Q

−5,000+50Q−5,000+50Q

Given Ajax’s pricing strategy, what is the profit-maximizing level of output for Ajax?

Ajax’s total dollar profit at this output level is.
Answer #1

MC =

MC = −5,000+50Q

Price (P) = $20,000
Profit maximizing level of output is that where P = MC
So, 20,000 = -5000 + 50Q
So, 50Q = 20,000 + 5,000 = 25,000
So, Q = 25,000/50
So, Q = 500

Total profit = Total Revenue(TR) - TC

TR = P*Q = 20,000*(500) = 10,000,000
TC=800,000−5,000Q+25Q2 = 800,000 - 5000(500) + 25(500)2 = 800,000 - 2,500,000 + 6,250,000 = 4,550,000

So, Total profit = 10,000,000 - 4,550,000 = $5,450,000